# Phonon Induced Contrast in Matter Wave Interferometer

### v1

@@ III Interactions in Stern-Gerlach Apparatus
where in different time intervals, the gradients take constant values ±$\pm\eta_{\text{b}}$. The closed loop is guaranteed as long as the acceleration and deceleration durations are the same.
[…]
Figure 2: Illustrations of matter wave packet trajectories shown in panel (a), velocities in panel (b) and acceleration protocols in panel (c), where blue lines represent CoM wave packet on the right ($S=S^{\text{R}}$) and orange dash lines represent the left ($S=S^{\text{L}}$). Duration of acceleration and deceleration in time intervals $[t_{1},~{}t_{2}]$, $[t_{2},~{}t_{3}]$, $[t_{4},~{}t_{5}]$, $[t_{5},~{}t_{6}]$ are the same and is denoted as a positive value $\tau_{a}$. During the time interval $[t_{3},~{}t_{4}]$, we assume a free-flight process of wave packets with zero acceleration. The duration of this process is denoted as $2\tau_{\text{f}}$. The total duration for wave packets flying in the SG apparatus is therefore $4\tau_{\text{a}}+2\tau_{\text{f}}$.
So far, we can observe from the Hamiltonian Eq. (7) that the diamond experiences two types of forces, i.e., the separation force acting on the NV-centre particle located at the centre $X_{c}$, and the diamagnetic force exerted by the entire diamond. Regarding the former, the force acts only on the central particle, the spin, which can be read out by doing partial derivative to the second term in Eq. (7), which has the form,
@@
where subscript represents the spin. From here, one can immediately write down the interaction between the magnetic field and phonon as $-F_{\text{s}}x_{c}$, which we will discuss in the following section. The tiny vibration of the NV-centre is imparted by $x_{c}(t)$, and the characteristic of the unique centre position is captured by the eigen-vector $Q^{c}_{q}(t)$ introduced in Eq. (2). Here, we keep labelling the spin-magnetic force as $F^{\text{L,R}}_{\text{s}}$ to distinguish it from the following diamagnetic force. But one should note that $F^{\text{L, R}}_{\text{s}}$ acts on the CoM, $X_{c}(t)$, and it is responsible for the trajectories of the diamond. The diamagnetic force acts on the entire diamond effectively at the positions of CoMs $X_{c}^{\text{L, R}}$. Here, we make an assumption that the diamagnetic force experienced by the entire diamond can be equally distributed to each individual atom inside the diamond.
$F_{i,\text{dia}}^{\text{L,R}}(t)=F_{c}^{\text{L,R}}(t)=\frac{\chi_{\rho}M}{N\mu_{0}}\left(B_{0}b+b^{2}X_{c}^{\text{L,R}}(t)\right).$ (11)
@@
Note that the diamagnetic force acting on each atom (including the NV centre for convenience) is equal to the diamagnetic force experienced by the CoMs of wave packets of diamond divided by the total number of atoms $N$. Here, $F_{c}^{\text{L,R}}$ represents the diamagnetic force experienced by any arbitrary atom $i^{\text{th}}$ inside the diamond. And the phonon-diamagnetic force coupling can now be written as $-F_{i,\text{dia}}^{\text{L,R}}x_{i}=-F_{c}^{\text{L,R}}x_{i}$ (summed over $i$ and mode $q$). Recalling that we use a linear magnetic field, this simple magnetic field is widely applied in interferometers Henkel et al. (2019); Marshman et al. (2022). Since the magnetic field gradient can be modelled as a step function, the separation force Eq. (10) acting on the spin is independent of its equilibrium position $X_{s(c)}$. Therefore, there is no coupling between the trajectories of wave packets and the phonon degrees of freedom. In contrast to the splitting force acting on the NV centre, the situation for the diamagnetic force (11) acting on each particle in the atomic chain is more complicated. The tiny vibrations of these particles (phonons) are coupled to their equilibrium positions (trajectories) through the magnetic field (considering $F_{c}^{\text{L,R}}\times x_{i}$ for each particle). Additionally, as shown in Fig. 1 (a), the forces acting on the NV centre in different wave packets are of the same magnitude but in opposite directions. Therefore, when the loop is completed, the contrast of phonons modes always differs by a positive or negative sign in both normal coordinates and momentum. For the diamagnetic force in Fig. 1 (b), the corresponding atoms in the atomic chain of the two wave packets experience the same direction of diamagnetic force but with different magnitudes (depending on the scale of spatial superposition). It is not difficult to imagine that the phonon modes will exhibit contrast reduction under such differential influences. In this toy model, we only investigate the contrast loss of phonon modes; the mismatch issue with the external degrees of freedom is beyond the scope of this work. We assume perfect overlap of the equilibrium positions (CoMs as well as $X_{i}^{\text{L,R}}$), a necessary assumption also employed in the seminal works Henkel and Folman (2022, 2023). Thus, we consider a simplified trajectory scenario where the trajectory of the diamond wave packets is solely governed by the spin-magnetic field coupling (10). In this simplest case, the equilibrium positions of corresponding particles in the two wave packets are symmetric given by $X^{\text{R}}=-X^{\text{L}}$, as illustrated in Fig. 2. The maximum accelerations of the two paths are thus a series of constants given by,
$a^{\text{L,R}}(t)=\frac{\mu S^{\text{L,R}}b(t)}{M}\,,$ (12)
we assume that at the initial moment, the spin inside the diamond has been prepared in a superposition, and the CoM of the diamond is at rest. When the loop starts, the magnetic field gradient $b(t)$ takes the protocol as mentioned in Eq.(9). In this toy model, we are neglecting the acceleration due to the induced diamagnetic potential for the time being. In fact, including the effect will not modify the current scenario a lot, as what matters is the relative force difference between the left and the right trajectories of the interferometer. As depicted in Fig. 2 (c), under the given magnetic field gradient settings, the acceleration of the CoMs also exhibits a series of step functions. The blue and orange line segments correspond to the trajectories of $S^{\text{R}}$ and $S^{\text{L}}$, respectively. The corresponding trajectories are shown in Fig. 2 (a), and they can be expressed by the following expressions,
$X_{c}^{\text{L,R}}(t)=\\ \left\{\begin{array}[]{lr}\frac{a^{\text{L,R}}\left(t+2\tau_{\text{a}}+\tau_{\text{f}}\right)^{2}}{2},&t=[t_{1},t_{2}]\\ \\ \frac{a^{\text{L,R}}\left(t^{2}-2\tau_{\text{a}}^{2}+2\tau_{\text{f}}t+\tau_{\text{f}}^{2}\right)}{2},&t=[t_{2},t_{3}]\\ \\ a^{\text{L,R}}\tau_{\text{a}}^{2},&t=[t_{3},t_{4}]\\ \\ \frac{a^{\text{L,R}}\left(t^{2}-2\tau_{\text{a}}^{2}-2\tau_{\text{f}}t+\tau_{\text{f}}^{2}\right)}{2},&t=[t_{4},t_{5}]\\ \\ \frac{a^{\text{L,R}}\left(-t+2\tau_{\text{a}}+\tau_{\text{f}}\right)^{2}}{2},&t=[t_{5},t_{6}]\end{array}\right.$ (13)
Thus, the relative separation distance $\Delta X(t)$ between the two trajectories is also a quadratic function of time. During the time interval $[t_{3},t_{4}]$, this distance between the two trajectories reaches its maximum value,
$\Delta X_{\text{m}}=2\left|a^{\text{L,R}}\right|\tau^{2}_{\text{a}},\$ (14)
where we defined the positive constant maximum acceleration $\left|a^{\text{L,R}}\right|=\mu\eta_{\text{b}}/M$, and total duration time $\Delta t=t_{6}-t_{1}$.
[…]
Figure 3: The occupation numbers of the phonons for the spin-magnetic field interaction shown in panel (a), and for the the diamagnetic force shown in panel (b). The blue and orange curves represent the wave packets of the diamond on the right and left sides, respectively. The black dashed line represents the number of phonons in the diamond before entering the SG device (scaled to one), and the blue and orange dashed lines represent the number of phonons on both sides after the closure of the trajectory of the diamond’s CoMs.
@@ IV Phonon fluctuations
In this section, we will investigate the impact of the above-mentioned interaction forces, see Eqs. (10, 11), acting on the lattice chain on the contrast of phonon wave packets. Here, we generally represent the aforementioned forces as $F_{i}(X_{i}^{\text{L,R}})$ (we will neglect $X_{i}^{\text{L,R}}$ for simplicity sometimes and use just $F_{i}^{\text{L,R}}$). The total Hamiltonian regarding every atom and their interactions with external fields can be written as,
@@
where $H_{\text{free}}$ is the free phonon Hamiltonian mentioned in Eq. (1). Eq. (15) represents the energy change of the diamond interacting with the magnetic field after entering the SG apparatus. When the diamond enters the SG apparatus, the interaction between its internal particles and the external field changes the number of phonons. With the Hamiltonian Eq. (15), the kinematic equations for the canonical coordinates of each vibrational mode $q$ can be derived as,
@@
Recall that $F_{i}^{\text{L,R}}$ in (16) is a generalized form introduced for convenience. When investigating spin-magnetic field coupling, it needs to be substituted with (10). In this case, the summation over atomic index $i$ disappears (we consider only a single NV-centre). When studying the influence of the diamagnetic force of the lattice chain in different trajectories on internal degrees of freedom, $F_{i}^{\text{L,R}}$ is replaced by (11), and the summation in (16) should include all atoms. In Fig. 3, we have separately illustrated the differences in the number of phonons induced by these two types of forces (1011). Recall that the eigenvector $Q_{i}^{q}$ of phonons is a quantity related to the atom index $i^{\text{th}}$. Here, we choose it to be one as the upper limit. the number of phonons is constant at the initial moment. After the diamond enters the SG apparatus, the number of phonons in the two trajectories changes differently and stabilizes with a constant difference after the duration time $\Delta t$. Recall that we are interested in diamond masses ranging from $M\sim 10^{-20}$ to $10^{-14}$ kg with frequencies $\omega_{q}$ in the range of $\omega_{q}\sim 10^{10}-10^{12}$ Hz. Therefore, the rate of change of the population of phonons of mode $q$ is extremely rapid. Here, for visual clarity, we plot the variation in the number of low-frequency phonons in two types of interactions. One should note that due to the size constraints of diamonds ($\omega_{0}=\pi c/L$), elastic waves in the material do not reach such low frequencies. The change of phonon number is small in the high-frequency range and becomes a perturbation. Here, we introduce the difference in phonon amplitude and momentum on both sides of the interferometer arms, which arise from $f_{q}^{\text{L,R}}$.
@@
where we define $\Delta f_{q}=f_{q}^{\text{R}}-f_{q}^{\text{L}}$. From here, one can observe that due to the spatial superposition of CoM’s, the lattice chain in the two wave packets experiences different forces. In Fig. 4, we illustrate the differences in the phase space of phonon wave packets from different paths. For visual clarity, the early stage of phase evolution is faded out, and normal coordinate and its conjugate have been made dimensionless. In Henkel and Folman (2022, 2023), the authors have shown that the phase space trajectories of phonons exhibit differences because the spin in the lattice is moving toward different directions, as shown in Fig. 4 (a). In the case of diamagnetic interactions, the differences in the force value exerted on the lattice still contribute to the reduction in contrast for phonon wave packets. In the next section, we will estimate the loss of contrast due to phonon fluctuations.
@@ V Loss of Contrast
From Eq. (20), it can be seen that the contrast is essentially the inner product of normalized phonon wave functions. Thus, a perfect closure of position, momentum, and other relevant degrees of freedom, including phonon vibrations, would yield $C_{q}=1$, while an imperfect closure results in $C_{q}\leq 1$. Contrast loss is defined as the logarithmic part of Eq. (20). We employ the time-dependent perturbation theory in quantum mechanics for the time evolution of phonon quantum states. In Eq. (15), the interaction between phonons and external field in the equation is considered as a small perturbation. Now, we ignore the summation over $q$ and focus on one mode; the Hamiltonian becomes,
@@
Here, recall that the force $F_{i}^{\text{L,R}}$ is time dependent and it is subjected to the equilibrium positions $X^{\text{L,R}}_{i}(t)$ of the atomic chain. Considering that the diamond is in a spatial superposition state, the forces acting on the atomic chains in the two trajectories are different. In the interaction picture, denote the time-dependent perturbation part in Eq. (21) as $V_{q}(t)$, the quantum state of phonon obeys the Schrödinger equation. The overlap of phonon wavepackets in the two arms becomes,
@@
[…]
Figure 4: Illustrations of phonon amplitude in x-axis and momentum in y-axis in phase space. Panel (a) represents the difference induced by spin-magnetic field interaction, and panel (b) shows the phase space due to phonon-diamagnetic force interaction. The blue and orange curves, respectively, represent the wave packets of the diamond on the right and left sides. The early stages of the phase space trajectories are faded for clarity.
In Eq. (23), one can first consider the case of spin-magnetic field coupling (10). In this case, $V_{q}^{\text{L}}$ and $V_{q}^{\text{R}}$ differ exactly by a positive or negative sign, which is because the forces exerted on the phonon wavepackets on both sides of the interferometer arms happen to be in opposite directions. In this scenario, the amplitudes and momentum of the phonons on both sides of the interferometer arms are displaced in opposite directions by $\frac{1}{2}\Delta u_{q}$ and $\frac{1}{2}\Delta\dot{u}_{q}$, respectively. The contrast of phonon can be comprehended as the probability of its amplitude and momentum staying in the initial state $\psi_{q}(0)$ Henkel and Folman (2022) as shown in Fig. 4 (a), once these two wave packets are separated, the overlap would not be perfect. For the diamagnetic interaction, the decrease in contrast is also not difficult to imagine. In this case, the amplitudes and corresponding momentum of the phonons on both sides of the interferometer arms are displaced to varying degrees in the same direction. One can always utilize Eq.(22) to calculate the contrast of the phonon wave function. (details are given in the appendix) as,
@@
Here, we have considered an initial thermal equilibrium ensemble of phonons composed of a state $\rho_{q}(0)$. Note that to evaluate the contrast Eq. (24), it is equivalent to evaluate the mean value of $U_{q}^{\text{L},\dagger}U_{q}^{\text{R}}$ which consists of both phonon’s “position” and its conjugate.Therefore, it is convenient to calculate the mean value by using the Wigner function, which is a quasi-distribution of phonon’s position and momentum in phase space, where for a harmonic mode considered here, it has a form of double Gaussian Eq. (6). Therefore, in the end, by multiplying the exponential operator with Eq. (6) and integrating over $u_{q}$ and $\dot{u}_{q}$, the contrast becomes,
@@
The final result of the contrast is an exponential function and is temperature-dependent. Naturally, when $\Delta\tilde{f}_{q}(\omega_{q})$ is zero, interference arms on both sides will exhibit no distinction, and Eq. (25) will yield a result of one. However, if there is a difference in the forces acting on the lattice chain on both sides of the interference arms, then $C\neq 1$, and the loss of contrast  [^4] From the expression of contrast, one can directly see why high temperature will destroy contrast. Suppose $T\rightarrow\infty$ for a fixed $\omega_{q}$ then the expression for the exponential part of Eq.(25) goes as, $-\frac{k_{\text{B}}T}{2\hbar^{2}\omega_{q}^{2}}|\Delta\tilde{f}_{q}|^{2}\rightarrow-\infty$, therefore $C\rightarrow 0$. The transfer function $\Delta\tilde{f}_{q}$ is always non-zero in our case. In the other limit, when $\omega_{q}\sim cq\rightarrow 0$, for a finite temperature, again means that the $C\rightarrow 0$. Lower values of $\omega_{q}$ can occur for systems with a lower speed of sound. Typically, these are more compressible objects for which the speed of sound is smaller than that of the diamond-like system. is expressed as $\log C$. The Fourier transform,
@@
Figure 5: We show the contrast reduction. Black and red lines correspond to spin-phonon and diamagnetic force-phonon interaction, respectively. Solid and long dash lines represent spin-magnetism and diamagnetic interaction in low temperature (4 K). Dot lines and short dash lines correspond to spin-magnetism and diamagnetic interaction in room temperature (300 K).
[…]
Figure 6: Panels (a), (b), (c), and (d) illustrate the contrast reduction resulting from spin-magnetic field coupling and diamagnetic interaction under different scenarios, where the colour shading from dark to light represents progressively intensified contrast reduction. The first column corresponds to spin-magnetic field coupling, while the second corresponds to diamagnetic interaction. In the first row, (a) and (b) depict scenarios with a diamond mass of $M\sim 2.25\times 10^{-14}$ kg, showing different maximum separations $\Delta X_{\text{m}}$ (horizontal axis) and duration times $\Delta t$ (vertical axis). In the second row, (c) and (d) assume a desired separation distance of $\Delta X_{\text{m}}=10^{-4}$m for the diamond’s CoMs. We explore the contrast loss caused by phonon wavepackets of different masses under varying duration times. In the four subfigures depicted, we have marked the cases where the contrast loss reaches $1\%$. In subfigures (a) and (b), for a diamond with mass $M=2.25\times 10^{-14}$kg, significant contrast loss requires achieving maximum separation distances of approximately $\sim 10^{-3}$m and $\sim 10^{-2}$m for the centre-of-mass wave packet within 0.05s, respectively. In subfigures (c) and (d), to achieve $\Delta X_{\text{m}}=10^{-4}$m within 0.05 seconds, the diamond mass needs to reach approximately $\sim 10^{-13}$kg and $10^{-12}$kg, respectively, to exhibit significant contrast loss. In these four cases marked, the magnetic gradients required are all large, becoming approximately $1000$ to $10000$ times greater than the current experimental capability.

### v2

@@ III Interactions in Stern-Gerlach Apparatus
where in different time intervals, the gradients take constant values ±$\pm\eta_{\text{b}}$. The closed loop is guaranteed as long as the acceleration and deceleration durations (denoted as $\tau_{\text{a}}$) are the same. In the interval $[t_{3},~{}t_{4}]$ there is a “free-flight” process with duration $\tau_{\text{f}}$, where the gradient is turned off, as illustrated in Fig. 2 (a).
In this paper, we concentrate on the contrast loss of phonon modes, while the mismatch issue in the CoM position and momentum of the two paths of the superposition is beyond the scope of this work. The maximum accelerations of the two paths are thus a series of constants given by,
$a^{\text{L,R}}(t)=\frac{\mu S^{\text{L,R}}b(t)}{M}\,,$ (11)
we assume that at the initial moment, the spin inside the diamond has been prepared in a superposition, and the CoM of the diamond is at rest. When the loop starts, wave packets on either sides of the interferometer behave distinctly according to $S^{\text{R}}$ and $S^{\text{L}}$, with opposite accelerations as illustrated in Fig. 2 (b). In this model, we neglect the acceleration due to the induced diamagnetic potential for the time being. In fact, including the effect will not modify the current scenario a lot, as what matters is the relative force difference between the left and the right trajectories of the interferometer.
In this simplest case, the CoM trajectories are symmetric given by $X_{\text{c}}^{\text{R}}=-X_{\text{c}}^{\text{L}}$, as depicted in Fig. 2 (d) they can be expressed by the following expressions,
$X_{c}^{\text{L,R}}(t)=\\ \left\{\begin{array}[]{lr}\frac{a^{\text{L,R}}\left(t+2\tau_{\text{a}}+\tau_{\text{f}}\right)^{2}}{2},{}{}&t=[t_{1},t_{2}]\\ \\ \frac{a^{\text{L,R}}\left(t^{2}-2\tau_{\text{a}}^{2}+2\tau_{\text{f}}t+\tau_{\text{f}}^{2}\right)}{2},&t=[t_{2},t_{3}]\\ \\ a^{\text{L,R}}\tau_{\text{a}}^{2},&t=[t_{3},t_{4}]\\ \\ \frac{a^{\text{L,R}}\left(t^{2}-2\tau_{\text{a}}^{2}-2\tau_{\text{f}}t+\tau_{\text{f}}^{2}\right)}{2},&t=[t_{4},t_{5}]\\ \\ \frac{a^{\text{L,R}}\left(-t+2\tau_{\text{a}}+\tau_{\text{f}}\right)^{2}}{2},{}{}&t=[t_{5},t_{6}],\end{array}\right.$ (12)
Thus, the relative separation distance $\Delta X(t)$ between the two trajectories is also a quadratic function of time. During the time interval $[t_{3},t_{4}]$, this distance between the two trajectories reaches the maximum value,
$\Delta X_{\text{m}}=2\left|a^{\text{L,R}}\right|\tau^{2}_{\text{a}},\$ (13)
where we define the positive constant maximum acceleration $\left|a^{\text{L,R}}\right|=\mu\eta_{\text{b}}/M$, and total duration time $\Delta t=t_{6}-t_{1}$.
(a) (a)
(b) (b)
(c) (c)
(d) (d)
Figure 2: The four sub-figures illustrate our magnetic field setting (a) inside the Stern-Gerlach apparatus and the corresponding accelerations (b), velocities (c) and trajectories (d) of CoM wave packets. In the time intervals $(t_{1},~{}t_{2})$ and $(t_{5},~{}t_{6})$ the magnetic gradient is $b(t)=+\eta_{\text{b}}$ while in $(t_{2},~{}t_{3})$ and $(t_{4},~{}t_{5})$ the gradient is switched to $b(t)=-\eta_{\text{b}}$. Due to the superposition of the spin ($S^{\text{L,R}}=\pm 1$), the CoM wave packets experience opposite splitting forces due to the NV. Thus in the left (red and light dash lines) and right (blue and solid lines) arms of interferometer CoMs possess opposite accelerations (donated by $a^{\text{L,R}}$), velocities and trajectories (denoted by $X^{\text{L,R}}$). The duration time for acceleration and deceleration are the same as $\tau_{\text{a}}$. We also assume a free-flight process between $t_{3}$ and $t_{4}$ with duration $2\tau_{\text{f}}$ where the magnetic gradient is turned off. Therefore, the entire time for the diamond flying inside the Stern-Gerlach apparatus is $4\tau_{\text{a}}+2\tau_{\text{f}}$.
III.1 Spin-Phonon Interaction
So far, we can observe from the Hamiltonian (8) that the diamond experiences two types of forces, i.e., the separation force acting on the NV particle located at $X_{\text{c}}$, and the diamagnetic force exerted by the entire diamond, which also effective located at CoM. Regarding the former, the force acts only on the central particle, the spin, which can be read out by doing partial derivative to the second term in Eq. (8). The splitting force has the form,
@@ III.1 Spin-Phonon Interaction
where the index “s” represents the spin. From here, one can immediately write down the spin-magnetic field-phonon interaction as $F_{\text{s}}x_{\text{s}}$ (recall that $x_{\text{s}}$ represent small vibration of the NV away from its equilibrium position). For clarity, in the following discussion relating to phonon, we will call this coupling as spin-phonon (Sp-Ph) interaction.
III.2 Diamagnetism-Phonon Interaction
For the diamagnetic force acting on the diamond, we consider that each particle has different magnetization opposite to the direction of the external magnetic field depending on its equilibrium position $X_{i}$. Therefore, each atom generates different magnitude of diamagnetic repulsion $F_{\text{dia}}^{\text{L,R}}(X_{i}):=F_{i,\text{dia}}^{\text{L,R}}$. From a macroscopic perspective, the overall diamagnetic repulsion experienced by the diamond is contributed by these atoms, thus we should obtain the following relationship (details of this part can been found in Appendix. A),
$F_{\text{c}}^{\text{L,R}}=\frac{\chi_{\rho}M}{\mu_{0}}\left(B_{0}b+b^{2}X_{\text{c}}^{\text{L,R}}\right)=\sum_{i}F_{i,\text{dia}}^{\text{L,R}}.$ (15)
where the second part of (15) comes directly from the third term of Eq. (8), representing the overall repulsion effectively exerted on the CoM. Therefore, the coupling between magnetic field (diamagnetism) and $i^{\text{th}}$ particle (phonon) in the chain can now be written as $F_{i,\text{dia}}^{\text{L,R}}x_{i}$. In the following discussion relating to phonon, we will call this coupling as diamagnetism-phonon (Dia-Ph) interaction.
Recalling that we used a linear magnetic field, this simple magnetic field is widely applied in interferometers Henkel et al. (2019); Marshman et al. (2022). Since the magnetic field gradient can be modelled as a step function, the separation force Eq. (14) acting on the spin is independent of its equilibrium position $X_{\text{s}}$. Therefore, there is no coupling between the trajectories of wave packets and the phonon degrees of freedom.
@@ III.2 Diamagnetism-Phonon Interaction
In contrast to the splitting force acting on the NV centre, the situation for the diamagnetic force (15) acting on each particle in the atomic chain is more complicated. The tiny vibrations of these particles (phonons) are coupled to their equilibrium positions (trajectories) through the magnetic field (considering $F_{i,\text{dia}}^{\text{L,R}}\times x_{i}$ for each particle). Additionally, as shown in Fig. 1 (a), the forces acting on the NV centre in different wave packets are of the same magnitude but in opposite directions. Therefore, when the loop is completed, the contrast of phonons modes always differs by a positive or negative sign in both normal coordinates and momentum. For the diamagnetic force in Fig. 1 (b), the corresponding atoms in the atomic chain of the two wave packets experience the same direction of diamagnetic force but with different magnitudes (depending on the scale of spatial superposition). It is not difficult to imagine that the phonon modes will exhibit contrast reduction under such differential influences.
(a) (a)
(b) (b)
Figure 3: Illustrations of occupation numbers of the phonons for spin-phonon interaction (panel (a)), and for the the diamagnetism-phonon interaction (shown in panel (b)). The blue and light-red dash curves represent the wave packets of the diamond on the right and left sides, respectively. The black dot line represents the number of phonons in the diamond before entering the SG device (scaled to one), and the blue and red dot lines represent the number of phonons on both sides after the closure of the trajectory of the diamond’s CoMs. Before the diamond enters the Stern-Gerlach interferometer, its wave packet is not split, so the initial number of phonons in the $q$ mode is determined only by temperature Eq. (5). After it enters the Stern-Gerlach interferometer, the number of phonons on both sides of the interferometer arms changes over time and a difference in quantity forms. The phonon frequencies shown here (2 MHz for (a) and 42 kHz for (b)) are very low, in order to achieve better visibility. For the diamond we are interested in (specifically $10^{-14}$ kg), its natural frequency reaches $10^{10}$ Hz. Similarly, at those frequencies, the number of phonons on both sides of the interferometer arms also forms a difference. However, since the frequency of the magnetic field variation (especially the $\sim 1$s required for the QGEM experiment) is much lower than the phonon frequencies, the change in the number of phonons in those frequency ranges is not drastic.
@@ IV Phonon fluctuations
In this section, we study phonon excited by the two forces (14, 15), acting on the lattice chain on the contrast of phonon wave packets. Here, we generally represent the aforementioned forces as $F_{i}(X_{i}^{\text{L,R}})$ (we will neglect $X_{i}^{\text{L,R}}$ for simplicity sometimes and use just $F_{i}^{\text{L,R}}$). The total Hamiltonian regarding every atom and their interactions with external fields can be written as,
@@
where $H_{\text{free}}$ is the free phonon Hamiltonian mentioned in Eq. (1). Eq. (16) represents the energy change of the diamond interacting with the magnetic field after entering the SG apparatus.
When the diamond enters the SG apparatus, the interaction between its internal particles and the external field changes the number of phonons. With the Hamiltonian Eq. (16), the kinematic equations for the canonical coordinates of each vibrational mode $q$ can be derived as,
@@
Recall that $F_{i}^{\text{L,R}}$ in (17) is a generalized form introduced for convenience. When investigating spin-phonon coupling, it needs to be substituted with (14). In this case, the summation over atomic index $i$ disappears (we consider only a single NV-centre). When studying the influence of the diamagnetic force of the lattice chain in different trajectories on internal degrees of freedom, $F_{i}^{\text{L,R}}$ is replaced by (15), and the summation in (17) should include all atoms.
In Fig. 3, we have separately illustrated the differences in the number of phonons induced by these two types of forces (1415). Recall that the eigenvector $Q_{i}^{q}$ of phonons is a quantity related to the atom index $i^{\text{th}}$. Here, we choose it to be one as the upper limit. the number of phonons is constant at the initial moment. After the diamond enters the SG apparatus, the number of phonons in the two trajectories changes differently and stabilizes with a constant difference after the duration time $\Delta t$. Recall that we are interested in diamond masses ranging from $M\sim 10^{-18}$ to $10^{-14}$ kg with frequencies $\omega_{q}$ in the range of $\omega_{q}\sim 10^{10}-10^{11}$ Hz. Therefore, the rate of change of the population of phonons of mode $q$ is extremely rapid. Here, for visual clarity, we plot the variation in the number of low-frequency phonons in two types of interactions. One should note that due to the size constraints of diamonds ($\omega_{0}=\pi c/L$), elastic waves in the material do not reach such low frequencies.
Here, we introduce the difference in phonon amplitude and momentum on both sides of the interferometer arms, which arise from $f_{q}^{\text{L,R}}$.
@@
where we define $\Delta f_{q}=f_{q}^{\text{R}}-f_{q}^{\text{L}}$. From here, one can observe that due to the spatial superposition of CoM’s, the lattice chain in the two wave packets experiences different forces. In Fig. 4, we illustrate the differences in the phase space of phonon wave packets from different paths. For visual clarity, the early stage of phase evolution is faded out, and normal coordinate and its conjugate have been made dimensionless. In Henkel and Folman (2022, 2023), the authors have shown that the phase space trajectories of phonons exhibit differences because the spin in the lattice is moving toward different directions, as shown in Fig. 4 (a). In the case of diamagnetism-phonon interaction, the differences in the force value exerted on the lattice still contribute to the reduction in contrast for phonon wave packets. In the next section, we will estimate the loss of contrast due to their couplings.
@@ V Loss of Contrast
From Eq. (21), it can be seen that the contrast is essentially the inner product of normalized phonon wave functions. Thus, a perfect closure of position, momentum, and other relevant degrees of freedom, including phonon vibrations, would yield $C_{q}=1$, while an imperfect closure results in $C_{q}\leq 1$ indicating the loss in contrast. We employ the time-dependent perturbation theory in quantum mechanics for the time evolution of phonon quantum states. In Eq. (16), the interaction between phonons and external field in the equation is considered as a small perturbation. Now, we ignore the summation over $q$ and focus on one mode; the Hamiltonian becomes,
@@
Here, recall that the force $F_{i}^{\text{L,R}}$ is time dependent and it is subjected to the equilibrium positions $X^{\text{L,R}}_{i}(t)$ of the atomic chain. Considering that the diamond is in a spatial superposition state, the forces acting on the atomic chains in the two trajectories are different. In the interaction picture, denote the time-dependent perturbation part in Eq. (22) as $V_{q}(t)$, the quantum state of phonon obeys the Schrödinger equation. The overlap of phonon modes in the two arms becomes,
@@
(a) (a)
(b) (b)
Figure 4: Splitting in the phase space of a fixed phonon mode where the horizontal and vertical axis represent phonon amplitude and momentum, respectively. Panel (a) represents the difference induced by Sp-Ph interaction, and panel (b) shows the differential phase due to Dia-Ph interaction. The blue and light-red dash curves, respectively, represent the wave packets of the diamond on the right and left sides. The early stages of the phase space trajectories are faded for clarity.
In Eq. (24), one can first consider the case of spin-phonon coupling (14). In this case, $V_{q}^{\text{L}}$ and $V_{q}^{\text{R}}$ differ exactly by a positive or negative sign, which is because the forces exerted on the phonon wave packets on both sides of the interferometer arms happen to be in opposite directions. In this scenario, the amplitudes and momentum of the phonons on both sides of the interferometer arms are displaced in opposite directions by $\frac{1}{2}\Delta u_{q}$ and $\frac{1}{2}\Delta\dot{u}_{q}$, respectively. The contrast of phonon can be comprehended as the probability of its amplitude and momentum staying in the initial state $\psi_{q}(0)$ Henkel and Folman (2022) as shown in Fig. 4 (a), once the two wave packets are separated, the overlap would not be perfect.
For the diamagnetism-phonon interaction, the decrease in contrast is also not difficult to imagine. In this case, the amplitudes and corresponding momentum of the phonons on both sides of the interferometer arms are displaced to varying degrees in the same direction. One can always utilize Eq.(23) to calculate the contrast of the phonon wave function. (details are given in the appendix) as,
@@
Here, we have considered an initial thermal equilibrium ensemble of phonons composed of a state $\rho_{q}(0)$.
Note that to evaluate the contrast Eq. (25), it is equivalent to evaluate the mean value of $U_{q}^{\text{L},\dagger}U_{q}^{\text{R}}$ which consists of both phonon’s “position” and its conjugate. Therefore, it is convenient to calculate the mean value by using the Wigner function, which is a quasi-distribution of phonon’s position and momentum in phase space, where for a harmonic mode considered here, it has a form of double Gaussian Eq. (7).
Therefore, in the end, by multiplying the exponential operator with Eq. (7) and integrating over $u_{q}$ and $\dot{u}_{q}$, the contrast becomes,
@@
The final result of the contrast is an exponential function and is temperature-dependent. Naturally, when $\Delta\tilde{f}_{q}(\omega_{q})$ is zero, interference arms on both sides will exhibit no distinction, and Eq. (26) will yield a result of one. However, if there is a difference in the forces acting on the lattice chain on both sides of the interference arms, then $C\neq 1$, and the loss of contrast  [^3] From the expression of contrast, one can directly see why high temperature will destroy contrast. Suppose $T\rightarrow\infty$ for a fixed $\omega_{q}$ then the expression for the exponential part of Eq.(26) goes as, $-\frac{k_{\text{B}}T}{2\hbar^{2}\omega_{q}^{2}}|\Delta\tilde{f}_{q}|^{2}\rightarrow-\infty$, therefore $C\rightarrow 0$. The transfer function $\Delta\tilde{f}_{q}$ is always non-zero in our case. In the other limit, when $\omega_{q}\sim cq\rightarrow 0$, for a finite temperature, again means that the $C\rightarrow 0$. Lower values of $\omega_{q}$ can occur for systems with a lower speed of sound. Typically, these are more compressible objects for which the speed of sound is smaller than that of the diamond-like system. is expressed as $\ln C$. The Fourier transform,